OpenStudy (anonymous):

y= -3x+5 5x - 4y = -3 solve by substitution. steps please. these are just some examples i need to see to get the hang of it.

OpenStudy (anonymous):

ok you know that \[y=-3x+5\] no matter what. so replace y in the second equation by \[-3x+5\] being careful with parentheses. you get \[5x-4(-3x+5)=-3\]

OpenStudy (anonymous):

5x - 4(-3x+5) = -3 5x +12x - 20 = -3 17x = 17 x = 1 plug x-1 in second equn. 5 - 4y = -3 4y = 8 y = 2 x=1,y=2

OpenStudy (anonymous):

then solve via jimmyrep steps. notice first that you need parentheses around \[-3x+5\] then you need to use the distributive law to get rid of the parentheses, i.e. distribute the -4

OpenStudy (anonymous):

that gets you to \[5x+12x-20=-3\] then combine like terms to get the next step \[17x-20=-3\] then add 20 to get \[17x=17\] then either think or divide by 17 to get \[x=1\]

OpenStudy (anonymous):

finally replace x by 1 in the first equation to find y \[y=-3(1)+5\] \[y=2\] and you answer is \[x=1,y=2\] or \[(1,2)\]

OpenStudy (anonymous):

thanks so much, i think i got it now. or i am pretty close to fully understanding. thanks for all your help